Hello out there,I'm having some problems that I need help with,the first one is,Express 0.0049 in scientific notation.The next one is,Express 7,200,000,000 in scientific notation.Next one is,A rectangle has side length 30 mm and diagonal length 34 mm.Find the length of the other side of the rectangle to the nearest whole number.And the last one is,A telephone pole is supported by a wire of length x.If the base of the wire x is 40 feet from the pole,which is 70 feet tall,what is the length of the wire to the nearest whole number.Thank you to whoever can help me.
Answer by checkley75(3666)

.0049=4.9*10^-3 we have moved the decimal 3 places to the right thus the power is 10^-3.
7,200,000=7.2*10^6 we have moved the 6 places to the left thus the power of 10 is 10^6.
33^2+x^2=34^2, 1089+x^2=1156, x62=1156-1089, x^2=67, x=sqrt67, x=8mm.
40^2+70^2=x^2, 1600+4900=x^2, x^2=6500, x=sqrt6500, x=81 feet for the wire.